# Large Language Model for Vulnerability Detection and Repair: Literature Review and Roadmap

### v1

@@ 5.1.3. Few-shot Prompting
Retrieval Augmentation. Retrieval augmentation is a technique used to enhance the few-shot prompting. It involves retrieving similar labeled data samples from the training set when given a test data sample and using these retrieved data as examples to guide the prediction of LLMs on the test sample. Liu et al. (Liu et al., 2023) proposed using efficient retrieval tools such as BM-25 and TF-IDF for retrieval. Zhou et al. (Zhou et al., 2024b) suggested using CodeBERT as a retrieval tool. This method first transforms code snippets into semantic vectors and then quantifies the similarity between two code snippets by calculating the Cosine similarity of their respective semantic vectors. It finally returned the top similar code based on the similarity scores.
5.2. How effective are the techniques used?
In this subsection, we examine the effectiveness of LLM adaptation techniques utilized in vulnerability detection. Specifically, we conduct a manual review of the ablation studies in the primary research to compare the performance of approaches with and without those LLM adaptation techniques. The differences observed can indicate the effectiveness of the LLM adaptation techniques. If a study does not provide such information, we exclude the results of the paper from this subsection.
Fig. 6 illustrates the effectiveness of LLM adaptation techniques in vulnerability detection. The blue dots represent the improvement rates reported in the primary studies, while the red triangle denotes the average improvement across the studies. When fine-tuning an LLM, domain-specific pre-training (DP), combination with program analysis (PA), combination with other DL modules (DL), data-centric innovations (DC), and causal learning (Causal) lead to improvements ranging from 3% to 16%, 4% to 16%, 1% to 8%, 2% to 17%, and 5%, respectively, in terms of Accuracy scores. For zero-shot prompting, we did not find a clear ablation study demonstrating the contribution of prompt engineering. For few-shot prompting, retrieval augmentation (RAG) has enhanced the effectiveness of LLMs, improving accuracy from 3% to 15%. These findings highlight the extent these methods were effective as reported in prior studies.
@@ 5.2. How effective are the techniques used?
Answer to RQ2: Our analysis revealed three commonly used techniques to adapt LLMs for vulnerability detection: including fine-tuning ($\approx$82%), zero-shot prompting ($\approx$11%), and few-shot prompting ($\approx$7%). We noted that different adaptation techniques led to varying improvements, with combination with program analysis (PA) technique achieving the highest average improvement.
@@ 6. RQ3: How are LLMs adapted for vulnerability repair?
Regarding the usage of LLMs, we summarize three major categories from the included studies: 1) fine-tuning (approximately 80%), which updates the parameters of LLMs using a labeled dataset, 2) zero-shot prompting (approximately 13%), which freezes the parameters of LLMs and does not consider any labeled data, and 3) few-shot prompting (approximately 7%), which also freezes the parameters of LLMs but considers a few labeled examples.
@@
In the following subsection, we will introduce detailed categories of LLM adaptation techniques for each of the following: 1) fine-tuning, 2) zero-shot prompting, and 3) few-shot prompting.
6.1. How are LLMs adapted for Vul. Repair?
6.1.1. Fine-tuning
@@
The fine-tuning process usually involves several steps/stages, such as data preparation, model design, and model training. Regarding fine-tuning, we categorize adaptation techniques into four groups based on the stages they mainly target: Diverse relevant inputs (data preparation), Model-centric innovations (model design), Domain-specific pre-training (model training), and Reinforcement learning (training optimization).
Figure 7. Adaptation Techniques of LLMs for Vul. Repair
@@
Diverse Relevant Inputs. Several studies have shown that in addition to the input vulnerable code, incorporating other diverse relevant inputs can boost the effectiveness of LLMs. Specifically, the AST of the vulnerable code input (Zhou et al., 2022), vulnerability descriptions (Zhou et al., 2022; Wei et al., 2023), and vulnerable code examples shared on CWE websites (Zhou et al., 2024a) can enhance the LLMs in repairing vulnerability. In addition, Wei et al. (Wei et al., 2023) found that vulnerability-inducing commits and vulnerability-fixing commits can also improve the effectiveness of the model. Lastly, Zhou et al. (Zhou et al., 2024a) found that certain vulnerable functions exceed the length limit of Transformer-based LLMs like CodeT5 (512 subtokens) (Zhou et al., 2024a). To address this limitation, they applied the Fusion-in-Decoder framework to partition a long code function into multiple segments and feed those segments into LLMs one by one. Their findings demonstrated that those segments can boost model effectiveness.
@@
Model-Centric Innovations. Model-centric innovations encompass methodologies that prioritize revising the model architecture of LLMs (i.e., the Transformer (Vaswani et al., 2017)). Specifically, Fu et al. (Fu et al., 2024) drew inspiration from Vision Transformer (VIT)-based object detection techniques in computer vision. They proposed vulnerability queries within the pre-trained Transformer model to identify vulnerable code blocks within the source code. Additionally, they trained a model to learn a vulnerability mask, enhancing the attention of the vulnerability queries on the vulnerable code areas.
@@
Domain-specific Pre-training. Domain-specific pre-training involves pre-training an LLM on data specific to a particular domain before fine-tuning it for the target task. Considering the similarity between bug-fixing and vulnerability-fixing tasks, several studies (Chen et al., 2023b; Chi et al., 2023; Zhang et al., 2023b; Zhou et al., 2024a) considered the task of bug-fixing as a pre-training task to enhance LLMs. Specifically, they first pre-trained LLMs on a bug fix corpus to fix bugs, and then fine-tuned LLMs on a vulnerability fix dataset to repair vulnerabilities. This kind of pre-training technique can be considered as Transfer Learning.
@@
Reinforcement Learning. Islam et al. (Islam and Najafirad, 2024) introduced SecureCode, an LLM tuned using a reinforcement learning framework. This approach integrated syntactic and semantic rewards to generate fixes for vulnerable code. Specifically, they leveraged the CodeBLEU (Ren et al., 2020) score as the syntactic reward and BERTScore (Zhang et al., 2020) as the semantic reward. After combining these rewards, they applied the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to fine-tune the CodeGen2-7B (Nijkamp et al., 2023b) model.
6.1.2. Zero-shot and Few-shot Prompting
For both zero-shot and few-shot prompting, we only observe one adaptation technique.
Prompt Engineering. For zero-shot prompting, Pearce et al. (Pearce et al., 2021) devised various prompts to guide LLMs in repairing vulnerabilities in a zero-shot manner. They incorporated error messages from static analysis tools like CodeQL or C sanitizers into their prompts, providing contextual information about the vulnerabilities. For instance, one prompt format is: ”// BUG: [vulnerability-description] //MESSAGE: [error-message] [Code] //FIXED VERSION:”, where [Code] represents the input vulnerable function. In addition, Wu et al. (Wu et al., 2023) also explored diverse prompts for directing LLMs in repairing vulnerabilities in a zero-shot manner. For example, they proposed a prompt that begins by commenting on the buggy code lines and adds ”BUG:” to highlight them. Then, the prompt appends ”FIXED:” at the end to guide the LLM in generating a corrected version of the code.
Furthermore, for few-shot prompting, Fu et al. (Fu et al., 2023) investigated a straightforward prompt strategy involving the provision of three repair examples within each prompt. This approach aimed to aid GPT-3.5 and GPT-4 in performing the repair task more effectively.
Figure 8. Effectiveness of Adaptation Techniques for Vulnerability Repair
6.2. How effective are the techniques used?
Fig. 8 illustrates the effectiveness of LLM adaptation techniques in vulnerability repair. The blue dots represent the improvement rates reported in the primary studies, while the red triangle denotes the average improvement across the studies. When fine-tuning an LLM, domain-specific pre-training (DP), combining LLM with diverse relevant inputs (DI), model-centric innovations (MC), and reinforcement learning (RL) have resulted in improvements ranging from 6% to 13%, 1% to 24%, 3% to 3.3%, and 6%, respectively, in terms of Accuracy scores. For zero-shot prompting and few-shot prompting, we did not find a clear ablation study demonstrating the contribution of prompt engineering. These results highlight the extent these methods were effective as reported in prior studies.
@@ 6.2. How effective are the techniques used?
Answer to RQ3: Our analysis revealed three commonly used techniques to adapt LLMs for vulnerability repair: including fine-tuning ($\approx$80%), zero-shot prompting ($\approx$13%), and few-shot prompting ($\approx$7%). We found that different adaptation techniques resulted in varying degrees of improvement, with domain-specific pre-training (DP) achieving the highest average improvement.

### v3

@@ 5.1. Fine-tuning
Retrieval augmentation is a technique used to enhance the few-shot prompting. It involves retrieving similar labeled data samples from the training set when given a test data sample and using these retrieved data as examples to guide the prediction of LLMs on the test sample. Liu et al. (DBLP:conf/dsc/LiuLGG23, ) proposed using efficient retrieval tools such as BM-25 and TF-IDF for retrieval. Zhou et al. (DBLP:journals/corr/abs-2401-15468, ) suggested using CodeBERT as a retrieval tool. This method first transforms code snippets into semantic vectors and then quantifies the similarity between two code snippets by calculating the Cosine similarity of their respective semantic vectors. It finally returned the top similar code based on the similarity scores. Du et al. (DBLP:journals/corr/abs-2406-11147, ) leverage a knowledge-level retrieval-augmentation framework consisting of three phases to identify vulnerabilities in a given code. First, they constructed a vulnerability knowledge base by extracting multi-dimensional knowledge from existing CVE instances using LLMs. Second, for a given code snippet, they retrieved relevant vulnerability knowledge from the constructed knowledge base based on functional semantics. Finally, they utilized LLMs to assess the vulnerability of the given code snippet by reasoning about the presence of vulnerability causes and proposing potential fixing solutions derived from the retrieved knowledge. Wen et al. (DBLP:journals/corr/abs-2404-15596, ) first retrieved the most relevant dependencies from call graphs based on a given input code, and then integrated these dependencies with the input code into LLMs to detect vulnerabilities.
@@
Answer to RQ2: Our analysis revealed three commonly used techniques to adapt LLMs for vulnerability detection: including fine-tuning ($\approx$73%), prompt engineering ($\approx$17%), and retrieval augmentation ($\approx$10%).
Figure 7. Adaptation Techniques of LLMs for Vul. Repair
@@ 6. RQ3: How Are LLMs Adapted for Vulnerability Repair?
Regarding the usage of LLMs, we summarize three major categories from the included studies: 1) fine-tuning (approximately 63%), which updates the parameters of LLMs using a labeled dataset, and 2) prompt engineering (approximately 37%), which freezes the parameters of LLMs and designs prompts to guide LLMs in generating relevant responses.
@@
In the following subsection, we will introduce detailed categories of LLM adaptation techniques for each of the following: 1) fine-tuning and 2) prompt engineering.
6.1. Fine-tuning
Some studies (DBLP:conf/sigsoft/FuTLNP22, ; DBLP:journals/ese/FuTLKNPG24, ) employ default fine-tuning methods to adapt LLMs for vulnerability repair. In this subsection, we focus on research that introduces advanced fine-tuning techniques that extend beyond the default approach to achieve improved effectiveness.
@@ 6.1. Fine-tuning
Regarding fine-tuning, we categorize advanced adaptation techniques into four groups based on the stages they mainly target: Data-centric innovations (data preparation), Model-centric innovations (model design), Domain-specific pre-training (model training), and Reinforcement learning (training optimization). Table 3 summarizes the existing LLM-based methods for vulnerability repair.
@@
Data-Centric Innovations. Several studies have shown that in addition to the input vulnerable code, incorporating other diverse relevant inputs can boost the effectiveness of LLMs. Specifically, the AST of the vulnerable code input (zhou2022spvf, ), vulnerability descriptions (zhou2022spvf, ; DBLP:journals/ejwcn/WeiBWLYSL23, ), and vulnerable code examples shared on CWE websites (zhou2024out, ) can enhance the LLMs in repairing vulnerability. In addition, Wei et al. (DBLP:journals/ejwcn/WeiBWLYSL23, ) found that vulnerability-inducing commits and vulnerability-fixing commits can also improve the effectiveness of the model. Lastly, Zhou et al. (zhou2024out, ) found that certain vulnerable functions exceed the length limit of Transformer-based LLMs like CodeT5 (512 subtokens) (zhou2024out, ). To address this limitation, they applied the Fusion-in-Decoder framework to partition a long code function into multiple segments and feed those segments into LLMs one by one. Their findings demonstrated that those segments can boost model effectiveness. Additionally, a study shows that removing less important parts of the code can help the model focus more effectively on the vulnerable lines, thereby improving performance. Berabi et al (DBLP:journals/corr/abs-2402-13291, ) presented a method for generating reduced code by leveraging static analysis outcomes, creating a compact representation of the program. This smaller version fits within the attention window of an LLM while preserving the crucial information required to learn a correct fix. The technique ensures that the reduced code snippet still produces the same static analysis report as the original program, maintaining the core problem for effective analysis.
@@
Model-Centric Innovations. Model-centric innovations encompass methodologies that prioritize revising the model architecture of LLMs (i.e., the Transformer (vaswani2017attention, )). Specifically, Fu et al. (fu2023vision, ) drew inspiration from Vision Transformer (VIT)-based object detection techniques in computer vision. They proposed vulnerability queries within the pre-trained Transformer model to identify vulnerable code blocks within the source code. Additionally, they trained a model to learn a vulnerability mask, enhancing the attention of the vulnerability queries on the vulnerable code areas. Islam et al. (DBLP:journals/corr/abs-2401-03374, ) proposed fine-tuning an LLM (CodeGen2 (DBLP:journals/corr/abs-2305-02309, )) to perform two tasks simultaneously. First, the model generates a fix for the vulnerable code, and second, it produces a developer-friendly description of the code as an explanation.
Table 3. Existing LLM-based Methods for Vulnerability Repair
LLMs LLM Adaptation Techniques Data Deployment
Size
Data-
centric
Model-
centric
Domain
Pre-training
Reinforce.
Learning
Default
Tuning
Prompt
Eng.
Input
Granularity
Real-world&
Tests
Interact
with Users
Integrated
to Workflow
Chen et al. (DBLP:journals/tse/ChenKM23, ) <1B Function
Chi et al. (DBLP:journals/tse/ChiQLZY23, ) <1B Function
Fu et al. (DBLP:conf/sigsoft/FuTLNP22, ) <1B Function
Zhou et al. (zhou2022spvf, ) <1B Function
Wu et al. (DBLP:conf/issta/WuJPLD0BS23, )
<1B,
1-10B,
>10B
Function
Fu et al. (fu2023chatgpt, ) >10B Function
Zhang et al. (zhang2023pre, ) <1B Function
Pearce et al. (pearce2023examining, )
<1B,
1-10B,
>10B
Function
Wei et al. (DBLP:journals/ejwcn/WeiBWLYSL23, ) <1B Function
Fu et al. (fu2023vision, ) <1B Function
Islam et al. (islam2024code, ) 1-10B Function
Fu et al. (DBLP:journals/ese/FuTLKNPG24, ) <1B Function
Zhou et al. (zhou2024out, ) <1B Function
Nong et al. (DBLP:journals/corr/abs-2402-17230, ) >1B Function
Berabi et al. (DBLP:journals/corr/abs-2402-13291, ) >1B Function
Ahmad et al. (DBLP:journals/corr/abs-2302-01215, ) >10B Function
Islam et al. (DBLP:journals/corr/abs-2401-03374, ) <10B Function
Nong et al. (nong2024automated, ) ¿10B Function
Wang et al. (DBLP:journals/corr/abs-2405-04994, ) ¿1B Function
@@
Domain-specific Pre-training. Domain-specific pre-training involves pre-training an LLM on data specific to a particular domain before fine-tuning it for the target task. Considering the similarity between bug-fixing and vulnerability-fixing tasks, several studies (DBLP:journals/tse/ChenKM23, ; DBLP:journals/tse/ChiQLZY23, ; zhang2023pre, ; zhou2024out, ) considered the task of bug-fixing as a pre-training task to enhance LLMs. Specifically, they first pre-trained LLMs on a bug fix corpus to fix bugs, and then fine-tuned LLMs on a vulnerability fix dataset to repair vulnerabilities. This kind of pre-training technique can be considered as Transfer Learning.
@@
Reinforcement Learning. Islam et al. (islam2024code, ) introduced SecureCode, an LLM tuned using a reinforcement learning framework. This approach integrated syntactic and semantic rewards to generate fixes for vulnerable code. Specifically, they leveraged the CodeBLEU (DBLP:journals/corr/abs-2009-10297, ) score as the syntactic reward and BERTScore (DBLP:conf/iclr/ZhangKWWA20, ) as the semantic reward. After combining these rewards, they applied the Proximal Policy Optimization (PPO) algorithm (schulman2017proximal, ) to fine-tune the CodeGen2-7B (nijkamp2023codegen2, ) model.
6.2. Prompt Engineering.
Zero-shot Prompting. For zero-shot prompting, researchers have sought to create effective prompts that guide LLMs in performing vulnerability repair. Prompt engineering typically includes components outlined below:
  • Vulnerability Descriptions. This component provides specific details about the types of vulnerabilities to be addressed, helping the LLM understand the nature of the issues within the code. Specifically, Pearce et al. (pearce2023examining, ) propose to use a format like “// BUG: [vulnerability-description]” to provide the descriptions on the vulnerability. Ahmad et al. (DBLP:journals/corr/abs-2302-01215, ) provided the bug description with explicit instructions for repair with LLMs, as illustrated in the prompt: “// BUG: Access Control Check Implemented After the Asset is Accessed. // Ensure that access is granted before data is accessed”.
  • Vulnerability Location. This element indicates the specific sections of the code where vulnerabilities may exist. Wu et al. (DBLP:conf/issta/WuJPLD0BS23, ) proposed a prompt that begins by commenting on the buggy code lines and adds ”BUG:” to highlight them, which provides the location of the vulnerability to LLMs.
  • Vulnerability-Related Auxiliary Information. This component encompasses auxiliary information about vulnerabilities. Nong et al. (DBLP:journals/corr/abs-2402-17230, ) incorporate vulnerability semantics—specifically, the behaviors of a vulnerable program that contribute to its susceptibility—into the prompt to enhance the performance of LLMs.
  • Program-Analysis Auxiliary Information. Wang et al. (DBLP:journals/corr/abs-2405-04994, ) proposed a program-analysis-based approach. Given a patch, their method first utilizes both the vulnerable code and its fixed version as input. The process begins by localizing the Minimum Edit Node (MEN), which is the common ancestor node of all modified nodes. Once the MEN is identified, customized rules are applied based on the MEN type to extract vulnerable code patterns. These patterns are then incorporated into the prompt to assist LLMs in generating more effective vulnerability fixes.
Few-shot Prompting. Furthermore, for few-shot prompting, Fu et al. (fu2023chatgpt, ) investigated a straightforward prompt strategy involving the provision of three repair examples within each prompt. This approach aimed to aid GPT-3.5 and GPT-4 in performing the repair task more effectively. Nong et al. (nong2024automated, ) proposed a novel few-shot prompting method. Given a vulnerable program with a known vulnerability location, they first narrowed the analysis scope to only the relevant subset of the program. Next, they elicited the LLM to identify the vulnerability’s root cause within this reduced scope, dynamically selecting exemplars from a pre-mined database that best fit the program. With the chosen exemplars, they formed the patching prompt to generate patches with LLMs.
@@
Answer to RQ3: Our analysis revealed three commonly used techniques to adapt LLMs for vulnerability repair: including fine-tuning ($\approx$63%) and prompt engineering ($\approx$37%).
7. RQ4: What are the characteristics of datasets and deployment?
A typical learning-based pipeline can be divided into three phases: 1) data preparation, 2) model construction, and 3) deployment. The previous RQs primarily focus on the model construction phase. In this RQ, we focus on examining the characteristics of the datasets used, as well as the deployment strategies employed in LLM-based vulnerability detection and repair studies.
7.1. Data Characteristics
In our investigation of the data utilized in LLM-based vulnerability detection and repair studies, we primarily focus on two aspects: 1) input code size and 2) data quality.
7.1.1. Input Granularity
Input granularity refers to the size or scope of the code provided to the model for vulnerability detection and repair. For example, in a vulnerability detection task with repo-level input granularity, the entire repository is given to the model, and it is tasked with identifying vulnerabilities within that repository. Input granularity can vary from the line-level, function-level, and class-level to repo-level.
For vulnerability detection, as shown in Table 2, the majority of studies concentrate on function-level vulnerability detection, with only 7 addressing line-level detection. The function-level approach analyzes entire functions for potential vulnerabilities, whereas line-level detection examines specific lines of code, which can be advantageous for pinpointing fine-grained locations of vulnerabilities. Regarding vulnerability repair, Table 3 reveals that all studies target function-level repairs.
However, there is a notable gap in research exploring larger input granularity for both vulnerability detection and repairs, such as class-level and repo-level approaches. Larger input granularity is essential for identifying vulnerabilities that span multiple functions. Specifically, there are no studies focused on class-level vulnerability detection and repair. For repo-level detection, one recent work (zhou2024comparison, ) initially explored the use of LLMs (e.g., Code Llama) to detect vulnerabilities across entire repositories, comparing their performance with SAST tools (e.g., CodeQL). Notably, none of the studied papers investigate repo-level vulnerability repair. Overall, research on vulnerability detection and repair for larger input granularity remains limited.
7.1.2. Data Quality
We separately investigate the data quality issues in vulnerability detection and repair. Specifically, we find that vulnerability detection studies primarily utilize datasets with labels generated by heuristic-based methods. Additionally, vulnerability repair studies often rely on datasets that lack test cases, making it difficult to fully evaluate the correctness of the generated repairs since there can be multiple correct fixes.
Vulnerability detection studies mainly rely on datasets using heuristic-based labeling methods. Table 2 shows that many studies have employed heuristic-based labeling functions to generate data for vulnerability detection tasks. Typically, they collect vulnerability-fixing commits from databases like NVD and label the pre-commit versions of the functions modified by these commits as vulnerable. Additionally, all unchanged functions in the same files are labeled as non-vulnerable. In this way, the vulnerable functions are those in the files affected by vulnerability-fixing commits. However, recent studies (DBLP:conf/icse/CroftBK23, ; DBLP:conf/issta/NieLWWLW23, ; DBLP:conf/raid/0001DACW23, ) have highlighted quality issues with vulnerability data generated by the heuristic-based method, including noisy or incorrect labels (e.g., labeling clean code as vulnerable).
Vulnerability repair studies mainly rely on datasets without test cases. As present in Table 3, obtaining real-world vulnerability data along with corresponding test cases is challenging. Most studies either utilize synthetic vulnerable code and the corresponding fixes (e.g.,  (islam2024code, )) or use real-world vulnerable code and fixes without test cases (e.g.,  (DBLP:journals/tse/ChenKM23, ; DBLP:journals/tse/ChiQLZY23, ; DBLP:conf/sigsoft/FuTLNP22, )). Two recent studies (DBLP:conf/issta/WuJPLD0BS23, ; pearce2023examining, ) utilized real-world vulnerability-fixing datasets with test cases, however, the number of samples is very limited, with only 42 and 12 samples being evaluated, respectively. This underscores the urgent need for more extensive real-world datasets and accompanying test cases.
7.2. Deployment Strategies
In our investigation of the deployment strategies utilized in LLM-based vulnerability detection and repair studies, we focus on two key aspects: 1) the communication and collaboration between developers and LLM-based methods, and 2) the integration of these methods into developers’ workflows. We focus on the communication and collaboration between users and LLM-based methods because they are essential in deploying LLMs in practice and are pivotal in maximizing the real-world value of LLMs (DBLP:journals/tkdd/YangJTHFJZYH24, ). Moreover, a recent survey (DBLP:conf/icse/Liang0M24, ) shows that developers desire opportunities for natural language interaction with LLM-based solutions, also highlighting the significance of this capability. In addition, seamless integration of these solutions into developers’ workflows is crucial for successful deployment. A recent study (DBLP:conf/seke/Zhang0ZAW23, ) found that even for the popular LLM-based tool GitHub Copilot, the top 1 limitation reported is the difficulty of integrating Copilot with IDEs or other plugins. This emphasizes the need for better support in integrating LLM-based tools with other frequently used tools in developers’ workflows such as IDEs.
Firstly, as shown in Table 2, limited studies on vulnerability detection in our review support communication or collaboration between developers and LLM-based methods. One exception is a study (yang2024security, ) where the method generates explanations for vulnerabilities during the detection process. However, interaction and communication with developers are essential; for instance, unexplained suggestions may be disregarded, as developers need to understand the rationale behind the recommendations to trust and implement them effectively. In vulnerability repair studies (as shown in Table 3), most also do not support communication between developers. An exception is a recent study by Islam et al. (DBLP:journals/corr/abs-2401-03374, ), which provides explanations for the fixes when addressing vulnerabilities.
Secondly, as illustrated in Table 2 and Table 3, all studies in our review are not integrated into developers’ workflows and tools but are evaluated only offline using static and historical data. For these solutions to be effectively adopted, they should seamlessly integrate into existing development environments and tools.
Answer to RQ4: The datasets used in LLM-based vulnerability detection and repair studies primarily focus on function-level or line-level inputs, with limited exploration of class or repository level. Many vulnerability detection datasets employ heuristic-based labeling methods, and repair datasets frequently lack associated test cases. Furthermore, current approaches do not emphasize integration with developer workflows or interaction and collaboration between developers and LLM-based models.